OpenStudy (selenagarcia889):

Use an equation to represent the corresponding angle relationship between A and B. Use the equation to find the measures of A and B. A = 4x and B = 2(x+14).

OpenStudy (selenagarcia889):

@Devonhoward15 @Lady.Liv1776

OpenStudy (faiqraees):

Corresponding angles equals to each other so A = B

OpenStudy (selenagarcia889):

Okay... Thanks but i really need help on the second part

OpenStudy (hurgleman):

I'm not 100% certain what you're actually looking for since it's been a while since I've worked with angle relationships. I don't remember how to find corresponding angles utilizing equations only... It's a mathematical system. What you need to do is solve for each letter. First, you want to put the equations vertically b = 2(x+14) a = 4x Now, solve one equation for x. b = 2(x+14) b/2 = x+14, b/2 - 14 = x a/4 = x Now you can substitute x in terms of A. b = 2(a/4 + 14) a = 4(b/2 - 14) Now, clean it up. b = a/2+14 a = 2b-56 From here, I'm not certain where to go, to be honest. I was trying to deal with them with systems using substitution, but I just kept getting equal equations.

OpenStudy (faiqraees):

A = B 4x = 2(x+14) 4x = 2x+28 2x = 28 x= 14 A = 4x = 4(14) = 56 B = 2(x+14) = 2(14+14) = 56